OpenStudy (anime_lover_levi):

(2sqrt(5)+3sqrt(7)^2 I know the answer is 21+2 sqrt(5) Whats the work i need to do to get there?

OpenStudy (anime_lover_levi):

@AloneS

OpenStudy (alones.):

Hold on.

OpenStudy (anime_lover_levi):

:3

OpenStudy (anime_lover_levi):

ok :)

OpenStudy (anime_lover_levi):

@LunyMoony

OpenStudy (lunymoony):

Solution 1 Rationalize the denominator: 3√7+5√√(7+5√√7+5√√)=3√7+5√√7+5√37+5(7+57+5)=37+57+5 3√7+5√−−−−−−√7+5√ 37+57+5 2 Rationalize the denominator: 3√7+5√√7+5√(7−5√7−5√)37+57+5(7−57−5) 3√7+5√−−−−−−√(7−5√)72−5√2 37+5(7−5)72−52 3 Simplify 7272 to 4949 3√7+5√−−−−−−√(7−5√)49−5√2 37+5(7−5)49−52 4 Apply this rule: x√2=xx2=x 3√7+5√−−−−−−√(7−5√)49−5 37+5(7−5)49−5 5 Simplify 49−549−5 to 4444 3√7+5√−−−−−−√(7−5√)44

OpenStudy (lunymoony):

This is cool, an online site that does square root problems http://www.cymath.com/answer.php?q=(sqrt(3))%2F(sqrt(7%2Bsqrt(5)))